OpenStudy (anonymous):

Solve the formula for 1/R1+1/R2=1/R3. Find R1

OpenStudy (anonymous):

(1/x) + (1/y) = (1/z) Find equation for x

OpenStudy (anonymous):

3x-x+2=4